# Preprint Manuscript

### v2

@@ 3 Construction
We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$. The construction proceeds in six stages: encode a proof sequence $\langle f_{1},\ldots,f_{N}\rangle$ using the $\beta$-function and Zeckendorf representation (3.1), implement Fibonacci addition for modus ponens verification (3.3), test axiom membership (3.3), enforce exactly-one justification per proof line (3.5), require the final line to be the target formula (3.6), and assemble the complete system (3.7).
@@ 3.4 Modus Ponens Verification
Procedure 3.9 (Greedy Witness Construction).
@@ 3.5 Proof Justification
Constraint 3.11 necessitates that exactly one of the following holds for each line $i$:
@@
Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 3.10), equation (3.56) necessitates:
@@
We consider the activation mechanism. If $b_{ax,i}=1$, then Constraint 3.5 forces $f_{i}\in\mathcal{A}$. If $b_{mp,i,j,k}=1$, then Constraint 3.10 forces $f_{i}=m_{i,j,k}=f_{j}+f_{k}$ (assuming the Fibonacci addition constraints hold). Constraint 3.11 ensures that exactly one of these activation variables is $1$ for each line $i$. This encodes the informal proof rule: "Each line is either an axiom or follows from two earlier lines by MP."
@@ 3.6 Target Formula Constraint
Constraint 3.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
@@ 3.7 Complete System Assembly
      By Constraint 3.3, each $f^{*}_{i}$ has a valid Zeckendorf representation.
@@
      By Constraint 3.11, each line $i$ is justified either as an axiom (Constraint 3.5) or by MP (Constraint 3.10).
@@
      Each $\varphi_{i}$ is either an axiom (see 3.3) or derived from earlier formulas by MP (by 3.12),
@@ 4.1 Merging via Sum-of-Squares
By the standard technique [Matiyasevich, 1970, 1993], the system
$\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$
with $\delta(P_{i})\leq 3$ is equivalent to:
$P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0.$ (4.63)
Since $\delta(P_{i})\leq 3$, we have $\delta(P_{\text{merged}})\leq 6$. Hence Exposition, $m=O(KN^{3})$ where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$.
@@ 4.3 Main Result
Corollary 4.3 (Cubic Incompleteness).
@@ 5.1 Degree Reduction Protocol
Procedure 5.2 (Monomial Shielding).
@@
Procedure 5.3 (Polynomial Degree Reduction).
@@ 5.3 Solvability and Complexity
For $P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ with $\delta(P_{i})\leq 3$ (from Table 1), we analyze the complexity of the polynomial reduction as follows. Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$. The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums. Since Section 3.7 establishes that $m_{0}=O(KN^{3})$ for the total number of constraints, Proposition 5.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$. Combining this with the $O(KN^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
@@ 6.2 Completeness
Step 3. Zeckendorf digits. For each $i\leq N$, compute the Zeckendorf representation using Procedure 3.9:
@@ 6.3 Undecidability Consequences
By Theorems 6.1 and 6.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$. By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$. Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 4.2), this exhibits a cubic equation whose solvability is undecidable. Encoding arbitrary Turing machine halting problems via similar constructions yields a uniform reduction from the halting problem to $\mathcal{D}_{3}$ solvability. ∎
@@ 6.4 Impredicativity
Thesis 6.5 (Impredicativity of Degree Threshold).
@@
Then there exists no first-order formula $B(k)$ in the language of arithmetic such that $T$ proves:
@@
Suppose such $B(k)$ exists with unique $k_{0}$ satisfying $\mathrm{T}\vdash B(k_{0})$. By the diagonal lemma, construct a Diophantine equation $E_{k_{0}}$ with $\delta(E_{k_{0}})=k_{0}$ whose solvability encodes $\neg B(k_{0})$ within $T$. If $E_{k_{0}}$ is solvable, then $T\vdash\neg B(k_{0})$, contradicting uniqueness. If $E_{k_{0}}$ is unsolvable, then $\mathcal{D}_{k_{0}}$ contains a decidable equation, contradicting clause (3). Hence no such $B(k)$ is definable in $\mathrm{T}$. ∎
@@
Thesis 6.6 (Diophantine Impredicativity).
@@
requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined. By Thesis 6.5, no closed term of this type exists in any consistent extension of $\mathsf{Q}$. The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
@@
    No uniform refutation exists (by Thesis 6.5, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
@@
Thesis 6.8 (Impredicativity of Degree Threshold).

### v3

@@ 4 Construction
We now construct a finite system of polynomial equations of degree at most $3$ such that the system has a solution over $\mathbb{N}$ if and only if the Fibonacci variant theory $T_{\text{Fib}}$ proves the target formula $G_{T}$.
@@ 4.4 Modus Ponens Verification
Procedure 4.9 (Greedy Witness Construction).
@@ 4.5 Proof Justification
Constraint 4.11 necessitates that exactly one of the following holds for each line $i$:
@@
Since $b_{ax,i},b_{mp,i,j,k}\in\{0,1\}$ (by 4.10), equation (4.69) necessitates:
@@
We consider the activation mechanism. If $b_{ax,i}=1$, then Constraint 4.5 forces $f_{i}\in\mathcal{A}$. If $b_{mp,i,j,k}=1$, then Constraint 4.10 forces $f_{i}=m_{i,j,k}=f_{j}+f_{k}$ (assuming the Fibonacci addition constraints hold). Constraint 4.11 ensures that exactly one of these activation variables is $1$ for each line $i$. This encodes the informal proof rule: "Each line is either an axiom or follows from two earlier lines by MP."
@@ 4.6 Target Formula Constraint
Constraint 4.13 forces $f_{N}=g_{T}$ (i.e., the final line of the proof is the target formula).
@@ 4.7 Complete System Assembly
      By Constraint 4.3, each $f^{*}_{i}$ has a valid Zeckendorf representation.
@@
      By Constraint 4.11, each line $i$ is justified either as an axiom (Constraint 4.5) or by MP (Constraint 4.10).
@@
      Each $\varphi_{i}$ is either an axiom (see 4.3) or derived from earlier formulas by MP (by 4.12),
@@ 5.1 Merging via Sum-of-Squares
We reduce the system $\Phi(\mathbf{x})$ from Theorem 4.16 to a single equation via sum-of-squares.
Degree bound. By Table 2, all polynomials $P_{i}\in\Phi$ satisfy $\delta(P_{i})\leq 3$ after guard-wrapping (Definition 3.6).
Merger. The system $\Phi=\{P_{1}=0,\ldots,P_{m}=0\}$ is equivalent to
$P_{\text{merged}}(\mathbf{x})=\sum_{i=1}^{m}P_{i}(\mathbf{x})^{2}=0,\quad\delta(P_{\text{merged}})\leq 6.$ (5.76)
Size. With $m=O(KN^{3})$ constraints (Remark Exposition) and $s_{i}=O(K)$ monomials per $P_{i}$, we obtain $O(K^{3}N^{3})$ total monomials.
@@ 5.3 Main Result
Corollary 5.3 (Cubic Incompleteness in $\mathbb{N}$).
@@ 6.1 Degree Reduction Protocol
Procedure 6.2 (Monomial Shielding).
@@
Procedure 6.3 (Polynomial Degree Reduction).
@@ 6.3 Solvability and Complexity
For $P_{\text{merged}}=\sum_{i=1}^{m_{0}}P_{i}^{2}$ with $\delta(P_{i})\leq 3$ (from Table 2), we analyze the complexity of the polynomial reduction as follows. Each squared polynomial $P_{i}^{2}$ has degree at most $6$ and contains at most $s_{i}^{2}$ monomials, where $s_{i}$ denotes the monomial count of $P_{i}$. The total monomial count is therefore $m=\sum_{i=1}^{m_{0}}s_{i}^{2}=O(m_{0}\bar{s}^{2})$, where $\bar{s}=\max_{i}s_{i}=O(K)$ by the bounds on Zeckendorf digit sums. Since Section 4.7 establishes that $m_{0}=O(K^{3}N^{3})$ for the total number of constraints, Proposition 6.6 implies that the number of auxiliary variables required is $O(m\log 6)=O(K^{3}N^{3})$. Combining this with the $O(K^{3}N^{3})$ base variables from Remark Exposition, the total variable count is $O(K^{3}N^{3})$.
@@ 7.2 Completeness
Step 3. Zeckendorf digits. For each $i\leq N$, compute the Zeckendorf representation using Procedure 4.9:
@@
7.3 Main Algorithm: Encoding Gödel’s Theorem
We present the complete reduction in algorithmic form:
[…]
Theorem 7.3 (Correctness of Algorithm 7.3).
The polynomial $P_{\text{final}}$ output by the algorithm satisfies:
$T\vdash G_{T}\quad\Longleftrightarrow\quad\exists\mathbf{x}\in\mathbb{N}^{m}:P_{\text{final}}(\mathbf{x})=0.$ (7.90)
Proof.
By Theorem 4.16 (soundness and completeness of Stages 1–6), Equation (5.76) (sum-of-squares equivalence), and Proposition 6.5 (degree reduction preserves solvability). ∎
@@ 7.4 Undecidability Consequences
By Theorems 7.1 and 7.2, solvability of $P_{\text{hard}}$ is equivalent to $\mathrm{T}_{\text{Fib}}\vdash G_{T}$. By [Gödel, 1931], if $\mathrm{T}_{\text{Fib}}$ is consistent, then $G_{T}$ is independent of $\mathrm{T}_{\text{Fib}}$. Since $\delta(P_{\text{hard}})\leq 3$ (Theorem 5.2), this exhibits a cubic equation whose solvability is undecidable. Encoding arbitrary Turing machine halting problems via similar constructions yields a uniform reduction from the halting problem to $\mathcal{D}_{3}$ solvability. ∎
@@ 7.5 Impredicativity
Thesis 7.6 (Impredicativity of Degree Threshold).
@@
Suppose such $B(k)$ exists with unique $k_{0}$ satisfying $\mathrm{T}\vdash B(k_{0})$. By the diagonal lemma, construct a Diophantine equation $E_{k_{0}}$ with $\delta(E_{k_{0}})=k_{0}$ whose solvability encodes $\neg B(k_{0})$ within $\mathrm{T}$. If $E_{k_{0}}$ is solvable, then $\mathrm{T}\vdash\neg B(k_{0})$, contradicting uniqueness. If $E_{k_{0}}$ is unsolvable, then $\mathcal{D}_{k_{0}}$ contains a decidable equation, contradicting clause (3). Hence no such $B(k)$ is definable in $\mathrm{T}$. ∎
@@
Thesis 7.7 (Diophantine Impredicativity).
@@
requires impredicative quantification over the decidability predicate $\mathsf{Decidable}(-)$, which itself depends on the threshold $k$ being defined. By Thesis 7.6, no closed term of this type exists in any consistent extension of $\mathsf{Q}$. The diagonal equation $E_{k_{0}}$ acts as a negative membership witness:
@@
    No uniform refutation exists (by Thesis 7.6, any such refutation enables constructing $E_{k_{0}}$, violating consistency).
@@
Thesis 7.9 (Impredicativity of Degree Threshold).